OpenStudy (toxicsugar22):

According to the following reaction, how many moles of water will be formed upon the complete reaction of 28.7 grams of hydrogen sulfide with excess oxygen gas? 2 H2S (g) + 3 O2 (g) = 2 H2O (l) + 2 SO2 (g)

OpenStudy (anonymous):

molar wt of H2S=2(1)+32=34 number of moles of sulphur=moles of water=34/28.7=1.184